please help with this
1. solve for x in 6x+12(10-x)=84
2. solve for d in -1/5(10d-5)=9
3.write an algebraic expression
"the product of a number and three is one half of five time the number decreased by one"

Answer by anjali_yamini(27)


1) 6x +12(10-x) = 84
6x + 120 - 12x = 84
-6x + 120 = 84
120 - 84 = 6x
36 = 6x
6 = x
check :
(6*6) +12(10-6)
= 36 +12*4
= 36 +48
= 84

2) -1/5(10d-5) = 9
-10d/5 +5/5 = 9
-2d +1 = 9
-2d = 9- 1 = 8
d = - (8/2)
d = -4
check :
-1/5(-40-5)
= 40/5 + 5/5
= 8 +1
= 9

3) let the number = n
so, 3 times the number = 3*n
= 3n
5 times the number = 5*n = 5n
so, 5 times the number minus 1 = 5n - 1
s0, half of 5 times the number - 1 = (5n-1)/2
so, 3n = (5n-1)/2